The product of 0.03 × 0.9 is: (a) 2.7 (b) 0.27 (c) 0.027 (d) 0.0027

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
0.03 × 0.9 can be written as = (3/100) × (9/10)

= 27/1000
On dividing a decimal by 1000, the decimal point is shifted to the left by three places
= 0.027

Thanks